OpenStudy (anonymous):

pleaseeeee help! The one-to-one functions g and h is defined as follows: g= {(-7,4),(-5,1),(-4,7),(7,-9)} h(x)=3x+8 find: g^-1(7)= h^-1(x)= (h o h^-1)(-7)=

OpenStudy (anonymous):

g={(-6,-2),(-2,-5),(7,5),(9,7)} h(x)=2x+9 find: g^-1(-2) = ? h^-1(x) = ? (h o h^-1)(7) = ?

OpenStudy (bahrom7893):

kumar please post your question in a new thread..

OpenStudy (anonymous):

?

OpenStudy (anonymous):

g^-1(x) is g inverse. When you have an inverse function, it means that the x and y values of your original function switch. So just switch all the x and y values for g(x) to get g^-1(x). This means g^-1(x) = {(-2,-6),(-5,-2),(5,7),(7,9)} From the points, we can see g^-1(-2) = -6. h^-1(x) is h inverse. So to figure out what the inverse function is, switch x and h(x) and solve for h(x), which becomes the inverse function. h(x) = 2x+9 becomes x = 2h(x)+9 x-9 = 2h(x) (x-9)/2 = h(x) Now because we solved for the inverse function, change h(x) to h^-1(x) = (x-9)/2 (h o h^-1)(7) is the same as h(h^-1(7)). From the previous problem, h^-1(x) = (x-9)/2, so we can substitute: h((7-9)/2) = h(-1) = 2(-1)+9 = 7

OpenStudy (anonymous):

thats not my question though... those are wrong numbers